THE CHART BELOW SHOWS THE NUMBER OF BOOKS READ IN A MONTH IN EACHOF MRS. GRAHAM'S LANGUAGE ARTS CLASSES.


1ST PERIOD 68 2ND PERIOD 72 3RD PERIOD 53 4TH PERIOD 77 5TH PERIOD 57
WHAT IS THE MEAN NUMBER OF BOOKS READ PER CLASS?
THE MEDIAN NUMBER OF BOOKS LIES IN WHAT CLASS PERIOD?
DOES A MODE EXIST FOR THIS SET OF DATA? WHY OR WHY NOT.

Answer by checkley71(8403)

(68+72+53+77+57)/5=327/5=65.4 IS THE MEAN (AVERAGE) NUMBER OF BOOKS READ BY THESE 5 CLASSES.
53,57,(68),72,77 THE MEDIAN IS THE MIDDLE NUMBER WHEN ARANGED IN ASSENDING ORDER & FALLS IN THE FIRST PERIOD.
THERE IS NO MODE BECAUSE NO NUMBER APPEARS MORE FREQUENTLY THAN ANY OTHER NUMBER.